find the residue of f(z) =1/(z2+1)2 at z=i


1
Expert's answer
2021-09-14T06:06:10-0400

"f(z) = \frac{1}{{{{\left( {{z^2} + 1} \right)}^2}}} = \frac{1}{{{{\left( {z + i} \right)}^2}{{\left( {z - i} \right)}^2}}}"

z = i is the zero of the order 2 denominator, so z = i is the second order pole.

Then

"\mathop {res}\limits_{z = i} f(z) = \mathop {\lim }\limits_{z \to i} {\left( {\frac{1}{{{{\left( {z + i} \right)}^2}{{\left( {z - i} \right)}^2}}} \cdot {{\left( {z - i} \right)}^2}} \right)^\prime } = \mathop {\lim }\limits_{z \to i} {\left( {\frac{1}{{{{\left( {z + i} \right)}^2}}}} \right)^\prime } = \mathop {\lim }\limits_{z \to i} \frac{{ - 2}}{{{{(z + i)}^3}}} = \frac{{ - 2}}{{{{(2i)}^3}}} = \frac{{ - 2}}{{ - 8i}} = \frac{1}{{4i}} = - \frac{i}{4}"

Answer: "\mathop {res}\limits_{z = i} f(z) = - \frac{i}{4}"